OpenStudy (anonymous):

The graph of y = ax2 + bx + c is a parabola that opens up and has a vertex at (0, 5). What is the solution set of the related equation 0 = ax2 + bx + c?

OpenStudy (anonymous):

@triston69 ı can solve

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

vertex ( r,k) = (0,5) a.(x-r)^2+k =y so a.x^2+5 =y = 0 a.x^2= -5 but parabol is open upward so a is not negative and x^2 is not negative so a.x^2 can not equal to -5 and no solution

OpenStudy (anonymous):



OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome